OpenStudy (anonymous):

please help me!! Solve for x and express roots in a + bi form: x^2/4 = 2x-10 and can you tell me how you got it?

OpenStudy (anonymous):

do you want to solve \[\frac{ x ^{2} }{4 }=2x-10\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x ^{2}=8x-40\] \[x ^{2}-8x=-40\] complete the squares by adding(co-efficient of x/2)^2

OpenStudy (anonymous):

i forgot how to complete the squares

OpenStudy (anonymous):

i told you add(-8/2)^2 ,i.e., 16

OpenStudy (anonymous):

ohhh thanks

OpenStudy (anonymous):

show me your work if any difficulty i will guide you.

OpenStudy (anonymous):

x^2 - 8x + 16 = -24

OpenStudy (anonymous):

write L.H.S as (a-b)^2 and R.H.S -1*24

OpenStudy (anonymous):

\[\left( x-4 \right)^{2}=4*6*-1=4*6*\iota ^{2}=\left( 2\sqrt{6}\iota \right)^{2}\]

OpenStudy (anonymous):

\[x-4=\pm 2\sqrt{6}\iota \] now write the two values of x

OpenStudy (anonymous):

x = \[4\pm 2i \sqrt{6}\]

OpenStudy (anonymous):

these are two solutions one with + sign and the other negative sign.

OpenStudy (anonymous):

\[x=4+2\sqrt{6}\iota ,and x=4-2\sqrt{6}\iota \]

OpenStudy (anonymous):

okay. i get it now. thank you so much!

OpenStudy (anonymous):

yw